# Thermodynamic speed limit for non-adiabatic work and its classical-quantum decomposition

### v1

@@ Appendix B Proof of the first main result
In this section, we show the first main result (14). To begin with, we use the diagonal-nondiagonal decomposition of the dissipator $\mathcal{D}[\rho_{t}]=\mathcal{D}_{\mathrm{d}}[\rho_{t}]+\mathcal{D}_{\mathrm{nd}}[\rho_{t}]$ introduced in Ref. [24], where
@@
being the “bath Hamiltnoian” that describes unitary part of the time-evolution induced by the bath [24].
@@
$=$ $\sqrt{4g_{tt}^{\rm QF}(\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}])},$
@@
$\left|\mathrm{Tr}[\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]\mathcal{D}_{\mathrm{d}}[\rho_{t}]]\right|\leq\sqrt{2A(\rho_{t})\dot{\sigma}},$ (54)
@@ B.1 Proof of (53)
We now derive inequality (53). The first line of (53) follows from the inequality [26, 30]
@@
where we use the Cauchy-Scwartz inequality and obtain the second line. We further note that
@@
$\langle n|\partial_{t}\rho_{t}|m\rangle=-\frac{i}{\hbar}(p_{m}-p_{n})\langle n|(H_{t}+H_{\mathcal{D}})|m\rangle=(p_{m}-p_{n})\langle n|\partial_{t}m\rangle,$ (60)
@@ B.2 Remarks on the quantum Fisher information metric
In this subsection, we give several remarks on the quantum Fisher information metric. Let us consider the Bures angle, which is the geodesic distance related to the quantum Fisher information metric, defined as [22, 65]
@@
is the Uhlmann fidelity. By evaluating the fidelity between $\rho_{t}$ and $\rho_{t}+d\rho_{t}$ and denoting $d\rho_{t}=\partial_{t}\rho_{t}dt$, we find up to the lowest order in $dt$ that [66]
@@ B.3 Proof of (54)
and a classical transition rate-like quantity [24]
@@
where the first line is obtained by using the Cauchy-Schwartz inequality, and the second line is obtained by using the relation $2(a-b)^{2}/(a+b)\leq(a-b)\ln\frac{b}{a}$ for $a,b\geq 0$. In Eq. (69), we note that
@@
We also use the following expression for the entropy production rate [24]
@@ Appendix C Proof of the second main result
$\mathcal{D}_{t}[\rho_{t}^{\rm diag}]=\sum_{n}|\epsilon_{n}\rangle\langle\epsilon_{n}|\mathcal{D}_{t}[\rho_{t}^{\rm diag}]|\epsilon_{n}\rangle\langle\epsilon_{n}|.$ (75)
@@
$|\dot{W}_{\rm cl}-\dot{W}_{\rm ad}|^{2}$ $=$ $|\mathrm{Tr}[\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]\mathcal{D}_{t}[\rho_{t}^{\rm diag}]]|^{2}$ (76)
$=$ $\Bigl{|}\sum_{n}\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle\langle\epsilon_{n}|\mathcal{D}_{t}[\rho_{t}^{\rm diag}]|\epsilon_{n}\rangle\Bigr{|}^{2}$
@@
$\leq$ $\sum_{n,\omega}\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle^{2}\gamma(\omega)\langle\epsilon_{n}|L_{\omega}\rho_{t}^{\rm diag}L_{\omega}^{\dagger}+\rho_{t}^{\rm diag}L_{\omega}^{\dagger}L_{\omega}|\epsilon_{n}\rangle$
@@
$q_{n}:=\langle\epsilon_{n}|\rho_{t}^{\rm diag}|\epsilon_{n}\rangle.$ (78)
@@
$=$ $\beta\sum_{\omega}\omega\gamma(\omega)\mathrm{Tr}[L_{\omega}^{\dagger}L_{\omega}\rho_{t}^{\rm diag}]+\sum_{\omega}\gamma(\omega)\mathrm{Tr}[L_{\omega}^{\dagger}L_{\omega}\rho_{t}^{\rm diag}\ln\rho_{t}^{\rm diag}-\ln\rho_{t}^{\rm diag}L_{\omega}\rho_{t}^{\rm diag}L_{\omega}^{\dagger}]$
$=$ $\beta\dot{Q}-\mathrm{Tr}[(\ln\rho_{t}^{\rm diag})\mathcal{L}_{t}[\rho_{t}^{\rm diag}]]$
@@
In the following, we show $\dot{\sigma}_{\rm cl}\leq\dot{\sigma}$ based on Ref. [40]. We first introduce a completely-positive and trace-preserving (CPTP) map
$\mathcal{E}^{\rm diag}_{t}[\chi]:=\sum_{n}|\epsilon_{n}(t)\rangle\langle\epsilon_{n}(t)|\chi|\epsilon_{n}(t)\rangle\langle\epsilon_{n}(t)|,$ (80)
which projects the density operator to diagonal states, i.e., $\rho_{t}^{\rm diag}=\mathcal{E}^{\rm diag}_{t}[\rho_{t}]$. We next introduce a CPTP map that describes the infinitesimal time-evolution $t\rightarrow t+\Delta t$ as
@@
and denote $\rho_{t+\Delta t}=\Lambda_{t,t+\Delta t}[\rho_{t}]$ and $\rho_{t+\Delta t}^{\rm diag}=\Lambda_{t,t+\Delta t}[\rho_{t}^{\rm diag}]$. It should be noted that $\Lambda_{t,t+\Delta t}[\mathcal{E}^{\rm diag}_{t}[\chi]]=\mathcal{E}^{\rm diag}_{t}[\Lambda_{t,t+\Delta t}[\chi]]$, and therefore $\rho_{t+\Delta t}^{\rm diag}=\mathcal{E}^{\rm diag}_{t}[\rho_{t+\Delta t}]$. We now express the difference between $\dot{\sigma}$ and $\dot{\sigma}_{\rm cl}$ as
$\dot{\sigma}-\dot{\sigma}_{\rm cl}$ $=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\Bigl{(}S(\rho_{t+\Delta t})-S(\rho_{t})-S(\rho^{\rm diag}_{t+\Delta t})+S(\rho^{\rm diag}_{t})\Bigr{)}$ (82)
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\Bigl{(}\mathrm{Tr}\Bigl{[}\rho_{t}\Bigl{(}\ln\rho_{t}-\ln\mathcal{E}^{\rm diag}_{t}[\rho_{t}]\Bigr{)}\Bigr{]}-\mathrm{Tr}\Bigl{[}\rho_{t+\Delta t}\Bigl{(}\ln\rho_{t+\Delta t}-\ln\mathcal{E}^{\rm diag}_{t}[\rho_{t+\Delta t}]\Bigr{)}\Bigr{]}\Bigr{)}$
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\Bigl{(}D(\rho_{t}||\mathcal{E}^{\rm diag}_{t}[\rho_{t}])-D(\Lambda_{t,t+\Delta t}[\rho_{t}]||\Lambda_{t,t+\Delta t}[\mathcal{E}^{\rm diag}_{t}[\rho_{t}]])\Bigr{)}\geq 0,$
where we used $\mathrm{Tr}[\rho_{t}\ln\mathcal{E}^{\rm diag}_{t}[\rho_{t}]]=\mathrm{Tr}[\rho_{t}^{\rm diag}\ln\rho_{t}^{\rm diag}]$ and obtained the second line, $D(\rho||\sigma):=\mathrm{Tr}[\rho(\ln\rho-\ln\sigma)]$ is the quantum relative entropy, and the last inequality follows from the monotonicity of the relative entropy [65].
@@ Appendix D Details of the example
In this section, we show additional details of the two-level system example discussed in Sec. VI. In the following, we use the notation $p_{\pm}=\langle\epsilon_{\pm}|\rho|\epsilon_{\pm}\rangle$ and $\rho_{+-}=\langle\epsilon_{+}|\rho|\epsilon_{-}\rangle$, and give explicit expressions of the non-adiabatic work and the bounds for two-level systems.
@@
  • The counter-diabatic Hamiltonian is given by
    and its variance reads
    $V_{\rho_{t}}(H_{\rm cd})=\dot{\theta}^{2}\left(1-4(\Im[\rho_{+-}])^{2}\right).$ (86)
@@
  • The activity-like quantity is given by
    $A(\rho_{t}^{\rm diag})=\frac{2\dot{\omega}_{t}^{2}A_{\rm act}}{(\gamma_{\downarrow}+\gamma_{\uparrow})^{2}},$ (88)
    where
    $A_{\rm act}=\gamma_{\uparrow}p_{-}+\gamma_{\downarrow}p_{+}$ (89)
    is the activity.
When deriving Eq. (88), we use the following explicit form of the the Drazin inverse [55] and its adjoint:

### v2

@@ Appendix B Proof of the first main result
In this section, we show the first main result (14). To begin with, we use the diagonal-nondiagonal decomposition of the dissipator $\mathcal{D}[\rho_{t}]=\mathcal{D}_{\mathrm{d}}[\rho_{t}]+\mathcal{D}_{\mathrm{nd}}[\rho_{t}]$ introduced in Ref. Funo et al. (2019), where
@@
being the “bath Hamiltonian” that describes unitary part of the time-evolution induced by the bath Funo et al. (2019).
@@
$=$ $\sqrt{4g_{tt}^{\rm QF}(\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}])}=:\mathcal{B}_{1}^{\rm nd},$
@@
$\left|\mathrm{Tr}[\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]\mathcal{D}_{\mathrm{d}}[\rho_{t}]]\right|\leq\sqrt{2A(\rho_{t})\dot{\sigma}}=:\mathcal{B}_{1}^{\rm d},$ (58)
@@ B.1 Proof of (57)
We now derive inequality (57). The first line of (57) follows from the inequality García-Pintos et al. (2022); Sekiguchi et al. (2024)
@@
where we use the Cauchy-Schwartz inequality and obtain the second line. We further note that
@@
$\langle n|\dot{\rho}_{t}|m\rangle=-\frac{i}{\hbar}(p_{m}-p_{n})\langle n|(H_{t}+H_{\mathcal{D}})|m\rangle=(p_{m}-p_{n})\langle n|\partial_{t}m\rangle,$ (64)
@@ B.2 Remarks on the quantum Fisher information metric
In this subsection, we give several remarks on the quantum Fisher information metric. Let us consider the Bures angle, which is the geodesic distance related to the quantum Fisher information metric, defined as Pires et al. (2016); Nielsen and Chuang (2010)
@@
is the Uhlmann fidelity. By evaluating the fidelity between $\rho_{t}$ and $\rho_{t}+d\rho_{t}$ and denoting $d\rho_{t}=\dot{\rho}_{t}dt$, we find up to the lowest order in $dt$ that Hübner (1992)
@@ B.3 Proof of (58)
and a classical transition rate-like quantity Funo et al. (2019)
@@
where the first line is obtained by using the Cauchy-Schwarz inequality, and the second line is obtained by using the relation $2(a-b)^{2}/(a+b)\leq(a-b)\ln\frac{b}{a}$ for $a,b\geq 0$. In Eq. (73), we note that
@@
We also use the following expression for the entropy production rate Funo et al. (2019)
@@ Appendix C Proof of the second main result
$\mathcal{D}_{t}[\rho_{t}^{\rm cl}]=\sum_{n}|\epsilon_{n}\rangle\langle\epsilon_{n}|\mathcal{D}_{t}[\rho_{t}^{\rm cl}]|\epsilon_{n}\rangle\langle\epsilon_{n}|.$ (79)
@@
$|\dot{W}_{\rm cl}-\dot{W}_{\rm ad}|^{2}$ $=$ $|\mathrm{Tr}[\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]\mathcal{D}_{t}[\rho_{t}^{\rm cl}]]|^{2}$ (80)
$=$ $\Bigl{|}\sum_{n}\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle\langle\epsilon_{n}|\mathcal{D}_{t}[\rho_{t}^{\rm cl}]|\epsilon_{n}\rangle\Bigr{|}^{2}$
@@
$\leq$ $\sum_{n,\omega}\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle^{2}\gamma(\omega)\langle\epsilon_{n}|L_{\omega}\rho_{t}^{\rm cl}L_{\omega}^{\dagger}+\rho_{t}^{\rm cl}L_{\omega}^{\dagger}L_{\omega}|\epsilon_{n}\rangle$
@@
$q_{n}:=\langle\epsilon_{n}|\rho_{t}^{\rm cl}|\epsilon_{n}\rangle.$ (82)
@@
$=$ $\beta\sum_{\omega}\omega\gamma(\omega)\mathrm{Tr}[L_{\omega}^{\dagger}L_{\omega}\rho_{t}^{\rm cl}]+\sum_{\omega}\gamma(\omega)\mathrm{Tr}[L_{\omega}^{\dagger}L_{\omega}\rho_{t}^{\rm cl}\ln\rho_{t}^{\rm cl}-\ln\rho_{t}^{\rm cl}L_{\omega}\rho_{t}^{\rm cl}L_{\omega}^{\dagger}]$
$=$ $\beta\dot{Q}-\mathrm{Tr}[(\ln\rho_{t}^{\rm cl})\mathcal{L}_{t}[\rho_{t}^{\rm cl}]]$
@@
In the following, we show $\dot{\sigma}_{\rm cl}\leq\dot{\sigma}$ based on Ref. Tajima and Funo (2021). We first introduce a completely-positive and trace-preserving (CPTP) map
$\mathcal{E}_{t}[\chi]:=\sum_{n}|\epsilon_{n}(t)\rangle\langle\epsilon_{n}(t)|\chi|\epsilon_{n}(t)\rangle\langle\epsilon_{n}(t)|,$ (84)
which projects the density operator to diagonal states, i.e., $\rho_{t}^{\rm cl}=\mathcal{E}_{t}[\rho_{t}]$. We next introduce a CPTP map that describes the infinitesimal time-evolution $t\rightarrow t+\Delta t$ as
@@
and denote $\rho_{t+\Delta t}=\Lambda_{t,t+\Delta t}[\rho_{t}]$ and $\rho_{t+\Delta t}^{\rm cl}=\Lambda_{t,t+\Delta t}[\rho_{t}^{\rm cl}]$. It should be noted that $\Lambda_{t,t+\Delta t}[\mathcal{E}_{t}[\chi]]=\mathcal{E}_{t}[\Lambda_{t,t+\Delta t}[\chi]]$, and therefore $\rho_{t+\Delta t}^{\rm cl}=\mathcal{E}_{t}[\rho_{t+\Delta t}]$. We now express the difference between $\dot{\sigma}$ and $\dot{\sigma}_{\rm cl}$ as
$\dot{\sigma}-\dot{\sigma}_{\rm cl}$ $=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\Bigl{(}S(\rho_{t+\Delta t})-S(\rho_{t})-S(\rho^{\rm cl}_{t+\Delta t})+S(\rho^{\rm cl}_{t})\Bigr{)}$ (86)
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\Bigl{(}\mathrm{Tr}\Bigl{[}\rho_{t}\Bigl{(}\ln\rho_{t}-\ln\mathcal{E}_{t}[\rho_{t}]\Bigr{)}\Bigr{]}-\mathrm{Tr}\Bigl{[}\rho_{t+\Delta t}\Bigl{(}\ln\rho_{t+\Delta t}-\ln\mathcal{E}_{t}[\rho_{t+\Delta t}]\Bigr{)}\Bigr{]}\Bigr{)}$
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\Bigl{(}D(\rho_{t}||\mathcal{E}_{t}[\rho_{t}])-D(\Lambda_{t,t+\Delta t}[\rho_{t}]||\Lambda_{t,t+\Delta t}[\mathcal{E}_{t}[\rho_{t}]])\Bigr{)}\geq 0,$
where we used $\mathrm{Tr}[\rho_{t}\ln\mathcal{E}_{t}[\rho_{t}]]=\mathrm{Tr}[\rho_{t}^{\rm cl}\ln\rho_{t}^{\rm cl}]$ and obtained the second line, $D(\rho||\sigma):=\mathrm{Tr}[\rho(\ln\rho-\ln\sigma)]$ is the quantum relative entropy, and the last inequality follows from the monotonicity of the relative entropy Nielsen and Chuang (2010).
@@ Appendix D Details of the example
In this section, we show additional details of the two-level system example discussed in Sec. VI. and give explicit expressions of the non-adiabatic work and the bounds for two-level systems.
@@
When deriving Eq. (42), we use the following explicit form of the the Drazin inverse Chen et al. (2021) and its adjoint: